OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 65 y + x = 7

OpenStudy (anonymous):

From the second equation, y = 7 - x So substitute 7 - x into x in the first equation...now work that out.

OpenStudy (anonymous):

(7 - x)^2 + x^2 = 65

OpenStudy (anonymous):

49 - 14x + x^2 + x^2 = 65

OpenStudy (anonymous):

2x^2 - 14x + 49 = 65

OpenStudy (anonymous):

2x^2 - 14x - 16 = 0

OpenStudy (anonymous):

2x^2-14x==16

OpenStudy (anonymous):

Factor out a 2: 2(x^2 - 7x - 8) = 0 x^2 - 7x - 8 = 0 (x - 8)(x + 1) = 0 x = 8 and x = -1 If x = 8, y = -1 If x = -1, then y = 8 Pairs of solutions are (8,-1) and (-1,8)

OpenStudy (anonymous):

(−8, 1) (1, 6) (6, 1) (9, −2)